OpenStudy (anonymous):

Which expression can be used to find the area of the circle below? 12 cm

OpenStudy (anonymous):



OpenStudy (kymber):

Weeelllll what are the options?

OpenStudy (cwrw238):

i'm assuming that 12 cm is the radius if so area = pi*12^2 = 144pi

OpenStudy (anonymous):

yes 12 is the radius

OpenStudy (cwrw238):

right so area = 144pi or 452.4 sq cms

OpenStudy (anonymous):

so Area = (3.14)(12^2)

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

thanks ill be a fan